1. Calculate the estimated calorie requirements for a 58-year-old female who is 5’8”, 160 lbs and is moderately active, using the Mifflin-St. Jeor Energy Estimation Formula.

2. Calculate the maximum amounts of saturated fat, monounsaturated fat, and polyunsaturated fat (in grams) that this person should consume.

3. Calculate the amount of total fat (in grams) that this person would require.

4. Calculate the amount of carbohydrates (in grams) that this person would require.

5. Calculate the amount of protein (in grams) this person would require, based on the Acceptable Macronutrient Distribution Range.

6. How many calories per day should this person consume to promote weight loss of one pound per week?

7. Taking into account the expenditure of 200 calories per day in physical activity,how many calories should this person consume daily to promote weight loss of two pounds per week?

8. Do you recommend the calorie level estimated in #7? Why or why not?
Answer #1

Answer :

1.ESTIMATED CALORIE REQUIREMENT:

Calorie requirement = BMR × 1.55

(1.55 is used for moderate activity).

Parameters given :

Weight = 120 lbs = 72.575 kg, height = 5'8" = 172.72 cms, Age = 58 yrs.

BMR = 10 × wt (in kg ) +6.25 × ht ( in cms) -5× age( in yrs)-161.

(161 is used for females).

BMR = 10 × 72.575 +6.25 × 172.72 -5×58-161

= 725.75+1079.5-290-161

= 1354.25

Estimated calorie intake = BMR ×1.55

=1354.25 ×1.55

=2099 kcal/day